### v1

# A backdoor attack against link prediction tasks with graph neural networks

Abstract
Graph Neural Networks (GNNs) are a class of deep learning models capable of processing graph-structured data, and they have demonstrated significant performance in a variety of real-world applications. Recent studies have found that GNN models are vulnerable to backdoor attacks. When specific patterns (called backdoor triggers, e.g., subgraphs, nodes, etc.) appear in the input data, the backdoor embedded in the GNN models is activated, which misclassifies the input data into the target class label specified by the attacker, whereas when there are no backdoor triggers in the input, the backdoor embedded in the GNN models is not activated, and the models work normally. Backdoor attacks usually occur when the training process of GNN models is not controlled, for example, backdoor attacks can occur when using models trained by third parties or using third-party data to train models. Backdoor attacks are highly stealthy and expose GNN models to serious security risks. Currently, research on backdoor attacks against GNNs mainly focus on tasks such as graph classification and node classification, and backdoor attacks against link prediction tasks are rarely studied. In this paper, we propose a backdoor attack against the link prediction tasks based on GNNs and reveal the existence of such security vulnerability in GNN models, which make the backdoored GNN models to incorrectly predict unlinked two nodes as having a link relationship when a trigger appear. The method uses a single node as the trigger and poison selected node pairs in the training graph, and then the backdoor will be embedded in the GNN models through the training process. In the inference stage, the backdoor in the GNN models can be activated by simply linking the trigger node to the two end nodes of the unlinked node pairs in the input data, causing the GNN models to produce incorrect link prediction results for the target node pairs. We demonstrate that this attack is effective and efficient by performing experimental evaluations on four popular models and four public datasets and comparing them with baseline, and the experimental results show that in the black-box scenario, this attack can achieve high attack success rates (more than 89 %) with small model accuracy losses (less than 1%) and small poisoning rates (about 1%).
1 Introduction
Graph-structured data is ubiquitous in the real world, covering various application domains such as social networks [1],urban transportation networks [2], knowledge graphs [3], and biological networks [4]. In order to efficiently process graph-structured data, various powerful graph neural network(GNN) models have been proposed. GNNs adopt the message passing mechanism to update the representations of nodes by aggregating information from their neighbors. GNNs can be used for various graph related tasks such as graph classification [5], node classification [6], link prediction [7, 8] etc. Link prediction aims to predict the existence of links between pairs of nodes in a graph, which is an important task of GNNs and has a wide range of applications such as friends recommendation in social networks [9], prediction of missing triples in knowledge graphs [10], and prediction of protein functions as well as interactions in biological networks [11].
GNNs have powerful performance in processing graph-structured data. However, recent studies [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22] have shown that GNNs are vulnerable to a new security threat called backdoor attacks. In backdoor attacks, the attacker injects specific patterns (called backdoor triggers) in the training data and then embeds the backdoor in the GNN models through training. In the inference phase, when the input samples contain the backdoor trigger, the backdoor in the models will be activated and perform the action specified by the attacker in advance. For example, the GNN models embedded with the backdoor misclassify input samples with triggers as attacker-specified target class label, whereas they perform normally when processing samples that do not contain backdoor triggers. Backdoor attacks often occur when the training process of GNN models are uncontrolled, e.g. backdoor attacks can occur by using models trained by third parties or by using third-party data to train the models [22]. Backdoor attacks are highly stealthy and they have posed serious security threat to GNNs.
Backdoor attacks have been extensively studied in computer vision and natural language processing [23, 24, 25, 26], while the threat of backdoor attacks on GNNs has not yet been fully explored. Current studies on backdoor attacks on GNNs have focused on the graph classification [12, 13, 14, 15, 16] and the node classification [16, 17, 18, 19], while there is few research on backdoor attacks against link prediction tasks.
To the best of our knowledge, the only existed backdoor attack studies on link prediction are LB [20] and DLB [21], where DLB is a backdoor attack method for link prediction on dynamic graphs, and LB is a backdoor attack method for link prediction on static graphs, which studies a similar problem to that of our work. Specifically, LB optimizes the structure and features of a random initial subgraph by using the gradient information of the target model, then obtains a dense subgraph as the trigger for the backdoor attack, and finally poisons the graph dataset with this subgraph, so as to embed the backdoor into the GNN models through training. Although LB achieves good results, this method requires a large attack budget because LB uses the gradient information of the target model to optimize the subgraph, and the dense subgraph makes the trigger easy to be detected.
In this paper, we investigate whether there are less costly and more stealthy method to implement backdoor attacks against link prediction. Specifically, we propose a backdoor attack method for link prediction with a single node as the trigger, which targets backdoor models that incorrectly predict link relationships between two unlinked nodes as linked when the trigger occurs. Initially, a new node is created as the trigger, and the features of the trigger node are generated by counting the frequency of occurrence of each feature in the original graph and selecting the subset of features with fewer occurrences; Then appropriate node pairs are selected in the training graph to inject that trigger for poisoning, and the backdoor is embedded in the model through the training process. In the inference phase, the implementation of the backdoor attack can activate the backdoor in the model by simply connecting the trigger node to the two end nodes of the target node pairs in the test data, causing the model to produce incorrect prediction results for that target node pairs. The illustration of our backdoor attack is shown in Figure 1. The left side of the figure shows the input graph and the red node is the backdoor trigger node. Part (a) of Figure 1 represents the input graph without the trigger, and node A and node B form a target node pair between them; Part (b) of Figure 1 represents the input graph with the trigger, where the red trigger node connect node A and node B of the target node pair respectively. When predicting the link relationship between nodes A and B with the model embedded with the backdoor, the backdoor model works correctly and correctly predicts that there is no link between node A and node B because there is no trigger present, while for the input of (b), the backdoor in the backdoor model is triggered due to the presence of the trigger and the backdoor model incorrectly predicts a link exists between node A and node B. We conduct experimental evaluations on four popular models and four benchmark datasets and compare them with baseline to demonstrate that the attack is effective and efficient, and the experimental results show that the attack success rate can reach more than 92 % in the black-box scenario with the poisoning rates of about 1%, and that the backdoor models accuracies loss less than 1%.
Our contributions in this paper are summarized as follows:
  1. 1.
    We propose a backdoor attack against link prediction with GNNs, revealing that GNNs are vulnerable to this security risk. The backdoor attack has following three features: the attack achieves better stealthiness by using only one single node as the trigger, and it requires lower poisoning rate in training dataset ; The attack can be implemented in black box scenarios, where the attackers do not need to know the parameters of target GNN models or create surrogate models, so it is easier or more practical for attacker to launch the attack; The attack has lower computing cost without involving any computation of gradients.
  2. 2.
    We propose a fast and effective trigger generation method, the core idea of which is to create the features of the trigger based on the statistical results of each element in the feature vector of each node in the graph, so as to increase the discrepancy between the trigger node and normal nodes to improve the effectiveness of the attack.
  3. 3.
    We propose a simple and effcient method for contaminating training data, which can achieve better attack performance with lower poisoning rates by connecting the trigger node to the selected nodes pairs in the graph.
  4. 4.
    We evaluate the performance of our attack with four popular models and four benchmark datasets and compare the results with baseline. The experimental results show that our attack can achieve more than 89 % attack success rates with poisoning rates of about 1 % in the black-box scenario, while the classification accuracy of the backdoored model on clean samples drops less than 1 % in comparison with that of the clean models.
The rest of the paper is organized as follows. In Section 2, we introduce the background knowledge of GNNs, link prediction and backdoor attacks; in Section 3, we introduce the adversarial attacks that GNNs are exposed to. Then we describe in detail the principle of backdoor attacks and the current state of study of backdoor attacks; in Section 4, we describe in detail our proposed method; in Section 5, we experimentally evaluate the performance of the proposed backdoor attack to demonstrate the effectiveness and feasibility of this backdoor attack; finally, Section 6 concludes the paper.
Figure 1: Illustration of our backdoor attack against link prediction. The trigger is the red node. Part (a) is the input graph without the trigger and node A and node B form a target node pair. The backdoor model works correctly and correctly predicts that there is no link between the node A and node B. Part (b) is the input graph obtained by adding the trigger from the graph in part (a) and the backdoor in the backdoor model is triggered due to the presence of the trigger and the backdoor model incorrectly predicts a link exists between the node A and node B.
2 Background
2.1 GNNs
Given an undirected and unweighted attribute graph $G=(V,E,X)$,where $V=\{v_{1},v_{2},\cdots,v_{N}\}$ is the set of nodes, $E$ is the set of edges, $X=R^{N\times d}$ represents the nodes attribute matrix, and the adjacency matrix $A\in R^{N\times N}$. For two nodes $v_{i},v_{j}\in V$, if $(v_{i},v_{j})\in E$, it means that there exists an edge between $v_{i}$ with $v_{j}$ and $A_{ij}=1$, otherwise, $A_{ij}=0$.$N=|V|$ is the total number of nodes, and $d$ is the number of node feature dimensions.
GNNs learn effective node representations $H^{k}\in R^{N\times F},k\in\{1,2,\cdots,K\}$ by combining graph structural information and nodes’ features through the message propagation mechanism, where $H^{k}$ indicates the representation of the k-th layer and $H^{0}=X$. Formally, we can define the framework of GNNs as follows:
$H^{k}=AGGREGATE^{k}\{A,H^{k-1},{\theta}^{k}\}$ (1)
where AGGREGATE function aggregates the neighboring features of each node in the graph to get new nodes’ features, $\theta$ is a learnable parameter. After obtaining the features of the nodes, they can be used for downstream tasks such as graph classification, node classification, link prediction, etc.
2.2 Link prediction
Link prediction is the prediction of potential links or missing links of node pairs in a network. Formally, given a graph $G=(V,E,X)$, all node pairs are represented by the set U in the graph $G$. $E_{u}=U-E$ denotes the links which are not present currently and there are potential links or missing links in $E_{u}$ in the graph. The purpose of link prediction is to predict these potential and missing links of node pairs. For clarity, in this paper, the node pair to be predicted for the existence or non-existence of link are all referred to as target node pair.
2.3 Link prediction methods based on GNNs
Link prediction methods based on GNNs have recently shown superior performance to traditional methods, which utilize nodes’ features and structural information in the graph to learn the embedding of the nodes. The advantage of GNNs is that they can adaptively capture complex patterns and features in the graph without relying on artificially designed heuristic rules or assumptions. In the case of Graph Auto-Encoder (GAE) [7], GAE computes the node representation $z_{i}$ for each node by using two layers of graph convolutional networks (GCNs) [6]
$\begin{split}Z=\tilde{D}^{-1/2}\tilde{A}\tilde{D}^{-1/2}\sigma&(\tilde{D}^{-1/2}\tilde{A}\tilde{D}^{-1/2}XW^{0})W^{1}\\ z_{i}&=Z_{i,:}\end{split}$ (2)
where Z is the node embedding matrix output by GAE, and the i-th row of Z is the embedding of node $z_{i}$. $\tilde{A}=A+I$ is the adjacency matrix of the self-connection of the given undirected graph $G$, $\tilde{D}=\sum_{j}A_{ij}$ is the degree matrix, $W^{0}$,$W^{1}$represent the weight matrices of the first and second layers of the GCNs respectively. $\sigma(\cdot)$is the RELU activation function. Then an inner product decoder is used to reconstruct the edges in the graph, that is to predict whether a link exists between two nodes. Specifically, for any pair of nodes $i$ and $j$, the probability that the link exists between them is:
$p_{\theta}(A_{ij}=1|X)=s(z_{i}^{T}z_{j})$ (3)
where $s(\cdot)$ is the sigmoid function. The objective function of the GAE is to maximize the log-likelihood of all edges (positive samples) and randomly sampling an equal number of non-edges (negative samples) as positive samples:
$L(\theta)=\sum_{(i,j)\in E^{+}}{logp_{\theta}(A_{ij}=1|X)}+\sum_{(i,j)\in E^{-}}{log(1-p_{\theta}(A_{ij}=1|X))}$ (4)
where $E^{+}$ is the set of edges and $E^{-}$ is the set of non-edges. GAE can be trained by stochastic gradient descent optimization algorithm.
2.4 Backdoor attacks
Backdoor attacks are a type of poisoning attack against deep learning models, which aims to implant some hidden malicious functions, i.e., backdoors, in the target models by using poisoned samples with triggers during the models training process. In the inference phase, for input samples that contain trigger, the backdoor will be activated, causing the target models to produce incorrect prediction results, e.g., the input samples are incorrectly misclassified to the target class label specified by the attacker, whereas the backdoor model works normally for benign samples that do not contain trigger. Backdoor attacks were first proposed in the field of images [23, 24]. Gu et al. [23] proposed the backdoor attack for the first time in the image domain, where the backdoor model is trained by poisoning some training samples using special markers as triggers, and in the inference phase, the backdoor model classifies the images with the triggers as the target label, and classifies the clean images normally. Subsequently, backdoor attacks in the textual domain have also been extensively studied [25, 26], e.g., Qi.et.al [26] used syntactic structures as the triggers for textual backdoor attacks.
Backdoor attacks are highly stealthy and harmful because, in the absence of triggers, the performance and behavior of backdoor models are not significantly different from normal models and are not easily detected and defended against. Backdoor attacks can occur in many scenarios, such as when using datasets, platforms, or models provided by third parties for training or deployment. Backdoor attacks pose a serious threat to the security and trustworthiness of deep learning systems, and require sufficient attention and concern.
3 Related works
In this section, we briefly introduce the adversarial and backdoor attacks on GNNs.
3.1 Adversarial attacks
Although GNNs have achieved excellent performance in a variety of graph learning tasks and have been used in many applications, recent studies have shown that GNNs are equally susceptible to adversarial attacks [27, 28, 29, 30, 31, 32, 33, 34, 35, 36]. Depending on the stage of data perturbation that occurs during the attack, adversarial attacks can be divided into evasion attacks and poisoning attacks. In evasion attacks, we assume that the target model is a given already trained model that cannot be modified by the attacker, and the attacker makes the model predictions wrong by adding perturbations to the test samples. Dai et al [27] proposed reinforcement learning, genetic algorithm, and gradient based methods to optimize the perturbation of graph structure to implement adversarial attacks respectively. Chen et al [28] implemented the attack by modifying the links to change the embedding of the nodes using a fast gradient approach. These studies [27, 28, 29, 30, 31] are practiced adversarial attacks by modifying the features of existing edges or nodes, which requires the attacker to have a high level of privilege, which is not practical in most of the cases in reality. Instead, some studies have explored more realistic approaches that do not modify the existing graph structure and node characteristics, but instead achieve this by injecting nodes and edges into the existing graph, which is more practical compared to the former. Wang et al [36] reduced the classification accuracy of GCN by injecting malicious nodes into the graph. In poisoning attacks, the attacker perturbs the training set before the models is trained, causing a severe degradation in the performance of the trained models. For example, NETTACK [29] initiates poisoning attacks by generating adversarial perturbations against node features and graph structures through the surrogate model; Metattack [32] initiates poisoning attacks by modifying the structure of the graph through meta-learning.
3.2 Backdoor attacks
In the graph data field, backdoor attacks have been less studied, and current works focuse on backdoor attacks against graph classification [12, 13, 14, 15, 16] and node classification [16, 17, 18, 19]. Xi et al. [12] proposed for the first time a backdoor attack method for GNNs that uses subgraphs as the triggers that can be dynamically customized for different graphs with different triggers to poison the data. Yang et al. [13] revealed a transferable graph backdoor attack without a fixed trigger pattern, which implements black-box attacks on GNN by the attacking surrogate models. Zheng et al [14] rethink the triggers from the perspective of motifs (motifs are frequent and statistically significant subgraphs of graphs containing rich structural information), proposed motif-based backdoor attacks, and gave some in-depth explanations of backdoor attacks. Xu et al.[16] select the optimal trigger position through interpretable methods to accomplish backdoor attacks against graph classification and node classification.
The above studies mainly focus on graph classification and node classification, and the only current work on backdoor attacks under the context of link prediction are LB [22] and DLB [21]. DLB investigated the link prediction backdoor attack on dynamic graphs, it generates different initial triggers by Generative Adversarial Network (GAN) [37], and then based on the gradient information of the attack discriminator in the GAN, selects some of the links of the initial triggers to form a trigger set to accomplish the backdoor attack. LB is similar to our work, but LB uses subgraphs as triggers and generates triggers by using the gradient information of the target models. Compared to his work, our proposed backdoor attack specifies the following different features: the use of a single node as the trigger, the low poisoning rate of the training samples, and more unnoticeable; the ability to implement a black-box backdoor attack without the need for the specific parameters of the target GNN models or constructing surrogate models, which makes the attack easier; and the lack of the need to use the gradient information, which reduces the amount of computation and improves the efficiency of the attack.
4 Method
In this section, we explain in detail how our attack is realized. Table 1 summarizes the the notions used in the following sections and their explanations.
Table 1: The explanations of the notions
Notation Explanations
$G=(V,E,X)$ Graph $G$ with edge set $E$, node set $V$ and attribute matrix $X$
$\lvert V\rvert$ The number of nodes
$f_{c}$ The clean link prediction GNN model
$n_{t}$ The trigger node
$x_{t}$ The features of the trigger node
$f_{b}$ The backdoored link prediction GNN model
$T$ Target link state
$E_{ij}$ The original-unlinked target node pair
$\widetilde{E_{ij}}$ The original-unlinked target node pair after embedding the trigger
$\alpha_{j}$ Frequency of occurrence of the i-th feature
$M_{k}$ the position indices of the nodes with the least frequent occurrence
$k$ The number of modified trigger features
$p$ The poisoning rate
$NPS$ The Target node pair scoring function
4.1 Attack overview
The backdoor attack proposed in this paper aims to the link prediction task under transductive learning, the goal of which is to incorrectly predict a link relationship between two unlinked target nodes. Formally, given a graph $G$ and a trigger $n_{t}$, the attacker generates the poisoned graph $\tilde{G}$ by embedding the trigger $n_{t}$ into the graph $G$. The backdoor model $f_{b}$ is obtained by training on $\tilde{G}$ and the clean model can be denoted as $f_{c}$. The backdoor attack target can be denoted as:
[…]
where $E_{ij}$ is the target node pair in graph G that are unlinked and $\widetilde{E_{ij}}$ denotes the target node pair after embedding the trigger. From equation 5, the first formula represents the effectiveness of the backdoor model, which means that the backdoor model $f_{b}$ predicts the link state of the original-unlinked target node pair $\widetilde{E_{ij}}$ embedded in the trigger to be the linked state $T$. The second formula represents the evasiveness of the backdoor model, which means that the backdoor model predicts the link state of the target node pair $E_{ij}$ without the trigger embedded with normal behavior and similar prediction accuracy as the clean model.
Figure 2: The proposed GNN backdoor attack framework. our backdoor attack consists of four steps: 1. The attacker gets the trigger node by analyzing the features of the nodes in the graph dataset. 2. The attacker selects the poisoned node pairs by a score function. 3. The attacker generates the poisoning graph by injecting the trigger into the above selected node pairs and change the link state between the above node pairs to be linked. And this dataset is used to train the GNN models to embed the backdoor in the models. 4. After the backdoor is embedded into the GNN models, the unlinked node pairs with the trigger node will activate the backdoor in the models, and they will be predicted as linked, while those without the trigger node will be predicted correctly.
Figure 2 depicts the process of this backdoor attack. It includes the following steps:
  1. 1.
    Trigger generation: the features of the trigger node are generated by counting the information of the features of all nodes in the training data and selecting the set of features indexes with the least frequency of occurrence, as detailed in Section 4.2 below;
  2. 2.
    Poisoned node pairs selection: a scoring function is proposed to select suitable node pairs by scoring each node pair in the training graph, which will be embedded in the trigger node for poisoning, as detailed in Section 4.3 below;
  3. 3.
    Backdoor injection: inject the trigger node into the poisoned node pairs selected in step 2 above (i.e., the trigger node is linked to two end nodes of the poisoned node pairs, respectively), and mark the linking relationship between the original-unlinked target node pairs as linked. The training graph dataset after injecting the trigger is called the poisoned training dataset. This poisoned training dataset is then used to train the GNN models, which learns those linked relationship between the trigger node and poisoned node pairs and injects it as the backdoor into the models, as detailed in Section 4.4 below;
  4. 4.
    Backdoor activation: in the inference phase, when the trigger node is connected to the end nodes of unlinked target node pair respectively, the backdoor models will incorrectly predict that there is a link between these two nodes. And when the trigger is not present, the backdoor models work normally.
The backdoor attack method proposed in this paper has the following assumptions:
  1. 1.
    This backdoor attack targets the link prediction under transductive learning. The goal of this backdoor attack is to incorrectly predict node pairs that are unlinked as linked.
  2. 2.
    The attacker can access to the training data, including the topology and features of the nodes, and the attacker can add additional edges and nodes to the graph to generate data for poisoning.
  3. 3.
    The attribute matrix of the graph is binary.
4.2 Trigger Generation
In this paper, we use a single node as the trigger and generate the features of the trigger node by the following two steps:
1.Determine the indexes of the position where ”1” occurs least frequently in the features of all nodes of the training dataset through statistics;
zheng et al. in motif-Backdoor [14] found that using non-existent or infrequent substructures as the triggers are more effective in implementing backdoor attacks. Inspired by this method, we use a single node as the trigger, select the position indexes in the dataset where the features of all nodes has the least frequency of ”1”, and then set the features value of the corresponding position indexes of the trigger’s features to ”1” according to the indexes of this position, which can explicitly separate the trigger node from the normal nodes in terms of the feature dimensions, so as to form a strong association between the trigger node and the target link state in the backdoor models obtained in the training process.
Specifically, in an undirected unweighted attribute graph $G=(V,E,X)$, where $V=\{v_{1},v_{2},\cdots,v_{N}\}$ is the set of nodes, $E$ is the set of edges, $X=R^{N\times d},j\in\{1,2,\cdots,d\}$ represents the nodes attribute matrix, and $X_{:,j}$ represents the individual column vectors in the node attribute matrix. In order to determine the position where ”1” occurs least frequently in the nodes attribute matrix, we count the number of 1st order paradigms of each column vector of the above matrix $X$ as follows:
$\alpha_{j}=\|X_{:,j}\|_{1},j\in\{1,2,\cdots,d\}$ (6)
Since we only consider the case where the attributes of the nodes in the graph are binary, the above formula calculates the frequency of occurrence of ”1” in each dimension of the features of each node, which is denoted as $\{\alpha_{1},\alpha_{2},\cdots,\alpha_{d}\}$. We select the top $k$ with the smallest values, denoted as $\{\alpha_{m_{1}},\alpha_{m_{2}},\cdots,\alpha_{m_{k}}\}$, where the index set $M_{k}=\{m_{1},m_{2},\cdots,m_{k}\}\subseteq\{1,2,c\dots,d\}$ stands for the k positions with the smallest frequency of ”1” in features of all nodes.
Considering that in general graph datasets such as Cora, CiteSeer, etc., the average number of features contained in nodes is generally about 1% of the total number of feature dimensions, and if the trigger node carries a large number of features can be easily detected, so in order to make our trigger node more unnoticable, we generally set $k$ to be 1% of the total number of feature dimensions.
2. Modify the features value of the corresponding positions of the trigger node to ”1” based on the position indexes above.
At first, all the elements of the initial features of the trigger node are set to ”0”, and then according to the above $k$ position indexes, the elements of the corresponding positions of the features of the trigger node are modified to ”1”. We denote the trigger node by $n_{t}$ and its i-th feature value by $x_{t}^{i}$, the above operation can be expressed as
$x_{t}^{i}=\left\{\begin{aligned} 1,\qquad i\in M_{k}\\ 0,\qquad i\notin M_{k}\\ \end{aligned}\right.$ (7)
After the above two steps, we have generated the trigger node, which will be used to poison the training dataset in the following.
4.3 Poisoned node pairs selection
After generating the trigger node features, the next step is to select the suitable node pairs to inject the trigger node to generate effective poisoned training data. Based on the message-passing mechanism of GNNs, the selection of poisoned node pairs is crucial for backdoor attacks.
In a graph, connected nodes may have the same features [38, 39]. A previous work [18] has shown that the embedding of the trigger destroys the node similarity in the original feature space, which will reduce the efectiveness of GNNs learning representations and decrease the efciency of downstream tasks. So the random injection of the trigger node destroys the node similarity of the feature space , and when the trigger node links different nodes, the GNNs aggregate confusing information that further misleads it to make wrong decisions[18]. To reduce this effect, we want all nodes embedded with the trigger to have links between them, i.e., the poisoned nodes form a complete graph between them, which can poison fewer nodes while the number of poisoned node pairs is fixed, and also ensures that all nodes containing trigger nodes will have links between them, which will enhance the effectiveness of the GNNs for backdoor feature learning. Besides, in order to be more conducive to the implementation of backdoor attacks, in the process of injecting the trigger node, we tend to select node pairs with sparse node features (meaning that the number of non-zero elements of the node’s features is as small as possible) to inject the trigger node. Because if our trigger node is connected to a node pair with dense features, then the backdoor model will have difficulty in distinguishing the features of the trigger node from the features of the normal nodes during the training process, which may lead to a decrease in the effect of the backdoor attack. In order to select node pairs with sparse features for posioning, we define the following node pairs scoring function (NPS) to measure the suitability of the node pairs in the training data to be used for injecting the trigger. The NPS function is defined as follows:
$NPS(v_{i},v_{j})=\|X_{v_{i}}\|_{1}+\|X_{v_{j}}\|_{1}$ (8)
where, $(v_{i},v_{j})$ is any unlinked node pair in the training data, $\|X_{v_{i}}\|_{1}$ and $\|X_{v_{j}}\|_{1}$ represent the number of non-zero elements of the feature vectors of nodes $v_{i}$ and $v_{j}$, respectively. As can be seen, the sparser the features of the node pairs, the smaller the NPS values.
We compute and sort all the unlinked node pairs in the training data according to the above NPS function, and select $q$ unlinked node pairs with the smallest NPS values as the target node pairs to be injected with the trigger, where, $q=p\cdot|V|$, $p$ is the poisoning rate, which represents the ratio of the number of poisoned node pairs to the total number of nodes in the graph.
4.4 Backdoor injection
After generating the trigger node and selecting the target node pairs for injecting the trigger, the next step is to inject the trigger node into the above selected poisoned node pairs and change the link state between them from unlinked to linked, thus obtaining the poisoned graph $\widetilde{G}$. Finally, by training the GNN models with the above poisoned training data, the backdoor is embedded in the target model. The overall algorithm flow is shown in algorithm 1.
4.5 Backdoor Activation
In the inference phase, when the trigger node is connected to two end nodes of unlinked target node pair respectively, the backdoor models incorrectly predict that there is a link between these two nodes. And when the trigger is not present, the backdoor models work normally.
Algorithm 1 Algorithm of our backdoor attack
0:  $G=(V,E,X)$,poisoning rate $p$ Calculate the occurrence frequency $\alpha_{j}$ of each feature by equation 6 Select the first k features with the lowest occurrence frequency to add to the index set $M_{k}$ Generate the trigger node $n_{t}$ by equation 7 Calculate the score of node pairs by equation 8 Select poisoned node pairs for poisoning by their scores Inject the trigger node into the poisoned node pairs Changes the link status of the poisoned node pairs to linked to generate poisoning graph $\widetilde{G}$ Train the model on poisoning graph $\widetilde{G}$ to obtain $f_{b}$
0:  backdoor model $f_{b}$ and trigger $n_{t}$
5 Experiment
To verify the effectiveness of the proposed backdoor attack method, we conduct the following experiments: initially, we test the attack success rates of the backdoor models generated under different poisoning rates on samples with the trigger and the prediction accuracy on clean samples to evaluate the attack effectiveness and stealthiness of backdoor models, as described in Section 5.2; secondly, we compare these results with the baseline to assess the effectiveness of the backdoor attack method proposed in this paper, as described in Section 5.3. Then we test the effects of different budgets $k$ on the attack success rates of the backdoor models and the prediction accuracy of the backdoor model on clean samples respectively, as described in Section 5.4. Finally we conducted the ablation studies to confirm the effectiveness of the scoring-based poisoned node pairs selection method used in the paper, as described in Section 5.5.
5.1 Experiment setting
5.1.1 Datasets
We evaluate our method in four graph benchmark datasets: Cora[40], CiteSeer[41], CS[42], and Physcics[42]. They are all undirected graphs composed of literature citation networks, where nodes represent papers, edges represent citation relations, node features are word vectors, and node labels are paper topics. Table 2 summarizes the basic statistics of these four graph databases.
Table 2: The dataset statistics
Datasets Nodes Edges Classed Features
Cora 2708 5429 7 1433
CiteSeer 3327 4732 6 3703
CS 18333 327576 15 6805
Physcics 34493 247862 5 8415
5.1.2 Link prediction models
In order to evaluate the effectiveness of our proposed attack method, we use four popular GNN models on link prediction, that is, Graph Auto-Encoder (GAE)[7], Variational Graph Auto-Encoder (VGAE) [7], Adversarial Regularized Graph Auto-Encoder (ARGA) [8], and Adversarial Regularized Variational Graph Auto-Encoder (ARVGA)[8]. We briefly describe these models as follows:
GAE[7]: GAE is the auto-encoder based graph embedding method that uses GCNs as an encoder to map the nodes in the graph to a low dimensional vector space and then uses inner product as a decoder to reconstruct the adjacency matrix of the graph.
VGAE[7]: VGAE is a graph embedding method based on variational auto-encoder, which differs from GAE in the encoder part. Instead of using GCNs directly to get the node vector representations, VGAE uses two GCNs to get the mean and variance of the node vector representations respectively, and then samples the node vector representations from a Gaussian distribution. This increases the diversity and robustness of the node vector representation.
ARGA[8]: ARGA is a graph embedding method based on adversarial regularization, which differs from GAE and VGAE in the regularization part.Instead of using KL scatter to constrain the node vector representations to obey a certain distribution, ARGA uses a discriminator to judge whether the node vector representations come from the real data distribution. This can make the node vector representation more consistent with the real data distribution, and at the same time avoid the loss of information caused by the KL scatter.
ARVGA[8]: ARVGA is also a graph embedding method based on adversarial regularization, which combines the techniques of variational auto-encoder and adversarial regularization. ARVGA employs two GCN networks to generate the mean and variance of the node vector representations, respectively, and uses a discriminator through adversarial training to improve the realism of the node representations.
5.1.3 Parameter setting
We construct GAE, VGAE, ARGA and ARVGA models for link prediction. For the parameters of the link prediction models, we retained the settings described in the corresponding papers. That is, For VGAE and GAE, we train for 200 iterations using Adam with a learning rate of 0.01. We use a 32-dim hidden layer and 16-dim latent variables in all. For ARGA and ARVGA, we train all auto-encoder related models for 200 iterations and optimize them with the Adam algorithm. Both learning rate and discriminator learning rate are set for the ARGA and ARVGA. rate and discriminator learning rate are set as 0.001. We construct encoders with a 32-neuron hidden layer and a 16-neuron embedding layer for all the experiments and all the discriminators are built with two hidden layers (16-neuron, 64-neuron respectively). In our experiments, we repeat the above steps five times and averaged the results of the experiments.
5.1.4 Metrics
To evaluate the effectiveness and evasiveness of the backdoor attack method proposed in this paper, we use the following evaluation metrics.
  1. 1.
    Attack success rate (ASR): Attack success rate (ASR) represents the percentage of node pairs that are unlinked embedded in the trigger that the backdoor model predicts to be in a linked state. It is as follows:
    $ASR=\frac{Number\ successful\ attack\ node\ pairs}{Total\ number\ of\ attack\ node\ pairs}$ (9)
  2. 2.
    Benign performance drop (BPD): To evaluate the evasiveness of our attack method,we utilize the area under the curve,AUC[43], to represent the accuracy of the link prediction model.If, in $n$ independent comparisons,the link that exists gets a higher score than the link that does not exist $n^{{}^{\prime}}$ times, and they get the same score $n^{"}$ times, then AUC is defined as
    $AUC=\frac{n^{{}^{\prime}}+n^{"}}{n}$ (10)
    We then use the Benign Performance Difference (BPD), which measures the difference in prediction accuracy between the clean GNN model and the backdoor GNN model on clean graph. The BPD is defined as
    $BPD=AUC_{c}-AUC_{b}$ (11)
    where $AUC_{c}$ and $AUC_{b}$ represent the accuracy of clean model and backdoor model respectively.Lower BPD represents better performance, which indicates that backdoor model and clean model are closer to each other in terms of accuracy on clean samples, and the backdoor attack have better stealth.
  3. 3.
    Poisoning rate (p):The poisoning rate $p$ represents the ratio of the number of target node pairs poisoned to the total number of nodes in the graph. The lower the poisoning rate, the easier and more stealthier the backdoor attack will be.
  4. 4.
    Budget(k): The budget $k$ respresents the number of modified trigger features. budget is defined as follows:
    $k=\lambda d$ (12)
    where $k$ is the budget, $d$ is the total number of node feature dimensions, and $\lambda$ represents the number of modified trigger node features as a percentage of the total number of node feature dimensions.
5.1.5 Baseline
In order to evaluate our attack method, we compare the only current paper on link prediction for static graph data, LB[22]as baseline, with our method.
5.2 The effects of different poisoning rates
We test the ASR and BPD of the backdoor models generated against the four target models GAE, VGAE, ARGA and ARVGA with different poisoning rates on four datasets to evaluate the impact of the poisoning rate on the effectiveness and evasiveness of the backdoor attack method proposed in this paper, where the poisoning rates in each dataset of the experiments are in order of 0.2%,0.5%,1%,2%, 5%. The results of the experiments are shown in the table 3.
Table 3: Impact of poisoning rate on ASR and BPD.
Datasets p ASR BPD
GAE VGAE ARGA ARVGA GAE VGAE ARGA ARVGA
Cora 0.2 93.66% 96.34% 90.88% 73.38% 0.11% -0.57% 0.45% -0.98%
0.5 98.40% 98.62% 96.66% 79.46% 0.52% -0.34% -0.01% 0.24%
1 99.53% 99.58% 97.66% 89.40% 0.02% -0.13% -0.33% -0.19%
2 99.06% 99.18% 98.26% 88.58% 0.27% 0.76% -0.19% 0.78%
5 99.62% 99.92% 99.16% 99.16% -0.18% 0.13% 0.07% 1.63%
CiteSeer 0.2 99.06% 98.94% 99.12% 96.48% 0.24% 0.46% -0.06% 0.43%
0.5 99.62% 99.90% 99.32% 98.72% -0.26% -0.70% -0.21% 0.19%
1 99.85% 99.84% 99.77% 99.63% 0.67% 0.64% 0.24% 1.08%
2 99.94% 99.98% 99.36% 99.58% 0.25% -0.15% 0.30% 0.08%
5 100.00% 99.98% 99.94% 99.93% 0.99% 0.12% 0.39% 0.11%
CS 0.2 95.08% 95.02% 96.08% 88.00% 0.56% 0.14% 0.02% -0.16%
0.5 97.48% 96.46% 97.56% 94.24% 0.04% -0.08% 0.02% -0.22%
1 97.89% 97.97% 98.22% 96.44% 0.04% -0.62% -0.06% 0.09%
2 98.62% 98.58% 98.64% 97.64% 0.15% 0.32% 0.12% 0.39%
5 99.10% 99.82% 98.92% 98.52% 0.17% 0.11% 0.06% -0.37%
Physics 0.2 98.78% 98.11% 93.95% 93.34% 0.36% -0.15% 0.07% 0.38%
0.5 99.32% 99.47% 97.73% 96.77% -0.61% 0.04% -0.04% -0.38%
1 99.59% 99.48% 98.62% 97.45% -0.25% 0.37% 0.02% -0.27%
2 99.79% 99.68% 99.76% 98.95% 0.25% 0.27% -0.11% 0.61%
5 100.00% 99.90% 100.00% 98.06% 0.22% 0.16% 0.17% 0.03%
Figure 3: The impact of the poisoning rate on ASR and BPD.
Figure 3 is a visualization of the above experimental results, in which the vertical coordinates of the top four plots are the ASR, the vertical coordinates of the bottom four plots are the BPD, and the horizontal coordinates are the poisoning rates. From Figure 3, we can observe that as p increases, the ASRs of the backdoor models increases over the four datasets. Taking the Cora dataset as an example, when p=0.2%, the attack success rates of the four backdoor models GAE, VGAE, ARGA and ARVGA are 93.66%, 96.34%, 90.88%,73.38% respectively, and their ASRs are increased to 99.53%, 99.58%, 97.66%,89.40% respectively when p=2%. Thus, our method requires only a small poisoning rate to achieve a good attack effect, which ensures that our attack is not easily detected.
From Figure 3, we can also observe that the BPD of the four backdoor models in the four datasets fluctuates somewhat with $p$, but is overall less than 1%. That indicates that the backdoor models generated by the attack method proposed in this paper have prediction accuracies for clean samples that are not much different from those of the clean models, which makes the backdoor models very stealthy.
5.3 Comparison with baseline
In this section we compare our backdoor attack method with baseline. We compare the attack performance of the backdoor models against baseline for the four link prediction models (GAE, VGAE, ARGA and ARVGA) on four datasets (Cora, CiteSeer, CS and Physcis), and then compare the attack performance of the backdoor model against Baseline for the GAE model on the large-scale dataset physics .
1.Comparison of attack performance against baseline on three datasets (Cora, CiteSeer, and CS).
We test the performance of our backdoor attack on three datasets (Cora, CiteSeer, and CS) for four link prediction models (GAE, VGAE, ARGA and ARVGA) and compared it with the experimental results of BASELINE as shown in table 4, where, the experimental results of baseline come from the original paper [22], whose experiments were conducted under the condition of 10% poisoning rate. Since our method requires only small poisoning rates to achieve high-precision attack effectiveness, we choose the experimental results with the 1% poisoning rate in table 4 to compare with the baseline.
Table 4: Comparison with the baseline.
Datasets models ASR BPD
baseline ours baseline ours
Cora GAE 81.25% 99.53% 2.82% 0.02%
VGAE 83.22% 99.58% 4.65% -0.13%
ARGA 92.56% 97.66% 1.08% -0.33%
ARVGA 85.25% 89.40% 2.49% -0.19%
CiteSeer GAE 82.25% 99.85% 3.13% 0.67%
VGAE 89.59% 99.84% 8.51% 0.64%
ARGA 98.78% 99.77% 4.97% 0.24%
ARVGA 81.55% 99.63% 1.60% 1.08%
CS GAE 50.61% 97.89% 1.13% 0.04%
VGAE 67.26% 97.97% 0.95% -0.62%
ARGA 63.52% 98.22% 0.32% -0.06%
ARVGA 68.67% 96.44% 2.86% 0.09%
From the experimental results in Table 4, the backdoor attack method proposed in this paper against the four target models outperforms baseline on all three datasets, achieving more than 89% ASRs on all of them. Taking the Cora dataset as an example, our method achieved 99.53%, 99.58%, 97.66%, and 89.40% ASRs on the four link prediction models based on GAE, VGAE, ARGA and ARVGA, respectively, which are higher than the baseline’s 81.25%, 83.22%, 92.56%, and 85.25%, respectively. All of our results on the Citeseer dataset achieve an ASR of more than 99%. On the CS dataset, the ASRs of baseline drops dramatically, while our method still achieves an over 96% ASRs at CS.
2. Comparison of backdoor attack performance with Baseline on the large-scale dataset
Considering that graph data in the real world are complex, Baseline specifically evaluates the effectiveness and evasiveness on the large-scale dataset physcics. Since baseline only tested the backdoor model for the GAE model with the poisoning rate of 10%, we chose the data related to the GAE backdoor model with the poisoning rate of 1% on the physcics dataset from Table 3 for comparison with it, as shown in the table 5.
From Table 5, we can see that baseline’s attack performance on large-scale datasets is lower than ours, and our proposed backdoor attack method still has high ASRs and evasiveness on the large-scale dataset.
Table 5: Comparison with the baseline on the large-scale dataset
Method ASR BPD
baseline 39.02% 0.78%
ours 99.86% 0.13%
5.4 Testing the impact of different budgets on ASR and BPD
In this section we evaluate the impact of different budgets $k$ on ASR and BPD. We test the effect of different budgets on our backdoor attack on four datasets (Cora, CiteSeer, CS, and Physics) for four link predictions models (GAE, VGAE, ARGA and ARVGA). In the experiments on each dataset,$\lambda$ takes values of 0.2%, 0.5%, 1%, 2%, and 5%, respectively. The experimental results are shown in table 6.
Table 6: The impact of budgets on ASR and BPD.
Datasets $\lambda$ ASR BPD
GAE VGAE ARGA ARVGA GAE VGAE ARGA ARVGA
Cora 0.2 96.28% 97.24% 82.57% 65.81% 0.47% -0.21% -0.04% 0.10%
0.5 99.21% 99.38% 92.41% 71.43% 0.07% -0.01% 0.17% 0.11%
1 99.53% 99.58% 97.66% 89.40% 0.02% -0.13% -0.33% -0.19%
2 99.68% 99.71% 99.04% 96.53% -0.62% $0$08% 0.46%
5 99.80% 99.89% 99.69% 99.14% 0.80% -0.17% -0.30% -0.85%
CiteSeer 0.2 99.38% 99.78% 97.17% 72.58% 0.72% -0.13% 0.21% 0.36%
0.5 99.95% 99.99% 99.53% 93.56% 0.96% 0.52% 0.01% 0.14%
1 99.85% 99.84% 99.77% 99.63% 0.67% 0.64% 0.24% 1.08%
2 99.95% 99.99% 99.85% 99.83% 0.02% -0.53% 0.70% -0.14%
5 100.00% 99.93% 99.89% 99.90% 0.60% 0.30% -0.78% -0.07%
CS 0.2 89.03% 87.76% 83.28% 70.86% -0.57% -0.18% $0$06%
0.5 94.96% 95.24% 95.39% 88.75% -0.20% 0.01% 0.12% -0.40%
1 97.89% 97.97% 98.22% 96.44% 0.04% -0.62% -0.06% 0.09%
2 98.10% 98.68% 99.02% 98.42% 0.14% 0.30% -0.01% 0.27%
5 98.53% 99.42% 99.49% 99.33% 0.73% -0.20% 0.01% 0.57%
Physics 0.2 95.89% 97.80% 86.21% 71.70% -0.56% -0.28% 0.02% 0.07%
0.5 98.74% 99.09% 96.31% 89.42% -0.72% -0.01% -0.16% 0.11%
1 99.59% 99.48% 98.62% 97.45% -0.25% 0.37% 0.02% -0.27%
2 100.00% 100.00% 99.53% 99.05% -0.33% -0.12% -0.05% -0.11%
5 99.92% 99.95% 99.87% 99.81% -0.41% 0.14% -0.06% 0.62%
Figure 4: The impact of budgets on ASR and BPD.
[…]
5.5 Ablation Studies
In order to verify the effect of the poisoning target node pairs selection method on the ASRs of the backdoor models,we conducted the following ablation studies. We generate poisoned training datasets by randomly selecting target node pairs to inject the trigger in the training sets of the four datasets (Cora, CiteSeer, CS and Physics), and retrain the four backdoor models and test their ASRs by using these poisoning training datasets respectively, and then compare the ASRs of the four backdoor models trained by selecting the target node pairs to inject the trigger through the NPS function as proposed in this paper, and the results of the experiments are shown in figeure 5. From the figure, we can see that the ASRs of random method is lower than that of ours. Especially on top of larger-scale graphs, such as CS and Physics, the ASRs of the Random method drop significantly, while our proposed NPS method still achieves excellent results. This is because our node pairs selection method chooses node pairs with sparse features, that are useful for our backdoor attack. The experimental results show that the poisoning method of injecting trigger by selecting target node pairs through the NPS function can effectively improve the ASRs of the backdoor attack.
Figure 5: Ablation studies.
6 Conclusion
This paper discusses the feasibility of implementing the backdoor attack in link prediction. Specifically, we implement a single-node injection backdoor attack strategy that exploits the differences in the features distribution of the dataset to select the optimal set of features for the trigger node and utilizes the features of the node pairs to find the suitlable node pairs for poisoning. Our empirical evaluation results on four representative link prediction models (GAE, VGAE, ARGA and ARVGA) and four benchmark datasets show that our backdoor attack achieves high ASRs with little impact on the GNNs accuracies of clean test graphs. Currently there is little research on graph backdoor defense, and in the future we will investigate defense methods against link prediction backdoor attacks.